OpenStudy (anonymous):

if a+b=3 and a^2+b^2=7, then what is the value of a^4 +b^4?

hartnn (hartnn):

can you find ab first ? from a+b and a^2+b^2 ?

hartnn (hartnn):

because, you'll need a^2b^2

OpenStudy (anonymous):

1+2=3?

OpenStudy (anonymous):

For the first part

hartnn (hartnn):

1+2 ? you guessing nu,bers ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

use this : \((a+b)^2 = a^2+b^2+2ab\) to find the value of ab can u?

OpenStudy (anonymous):

no quite

hartnn (hartnn):

you know values of a+b and a^2+b^2 just plug those in that formula...

hartnn (hartnn):

\((a+b)^2 = a^2+b^2+2ab\) (3)^2 = 7+2ab

hartnn (hartnn):

now ?

OpenStudy (anonymous):

Yes, thanks

OpenStudy (anonymous):

3/2=ab

hartnn (hartnn):

good, so once you have ab, use this (a^2+ b^2) ^2 = (a^4+ b^4) + 2 a^2b^2 you know a^2+b^2 , a^2b^2 , and you need a^4+b^4

hartnn (hartnn):

and ab is NOT 3/2 try again ?

OpenStudy (anonymous):

ab=1?

hartnn (hartnn):

yup, ab =1 is correct :)

OpenStudy (anonymous):

-1=a^4+b^4

hartnn (hartnn):

lets check (a^2+ b^2) ^2 = (a^4+ b^4) + 2 a^2b^2 7^2 = (....) + 2 (1) how you got -1 ? :O

OpenStudy (anonymous):

I got confused (a^2... with a+b

OpenStudy (anonymous):

47=a^4+b^4

hartnn (hartnn):

i get 47 too :)

OpenStudy (anonymous):

Thanks so much! You are awesome:)

hartnn (hartnn):

your welcome ^_^ you are a good learner too.